Sheepskin Company had the following transactions during 2006 Sales of $4,500 on accoun Collected $2,000 for services to be performed in 2007 Paid $625 cash in salaries Purchased airline tickets for $250 in December for a trip to take place in 2007 . 10. What is Sheepskins 2006 net income using cash basis accounting? A) $1,375 B) $5,625 C) $1,125 D) $5,875 11. What is Sheepskins 2006 net income using accrual accounting? A) S5,625 B) $3,875 C) $3,625 D) $5,875 12. Javiers Tune-Up Shop follows the revenue recognition principle. Javier receives a car on August 1, and services the car on September 1. The customer picks up the vehicle on ails the payment to Javier on September 15. Javier receives the check in the r 16. When should Javier show that the revenue was earned? September 2 and m mail on Septem A) September 17 B) September 16 C) September1 be D) September 15 E) August 1 13. Closing entries
media%2F849%2F849772dc-5c63-4a0c-90fe-38
Answer #1

10) Net income cash basis = 2000-625-250 = 1125

So answer is c) $1125

11) Net income accural basis = 4500-625 = 3875

So answer is b) $3875

12) Revenue recognize on september 1

So answer is c) Sep 1

13) Closing entries :

so answer is d) Cause the revenue and expense accounts to have zero balance